OpenStudy (jungleman03):

Rewrite f(x) = −(x + 3)2 − 1 from vertex form to standard form. f(x) = −x2− 6x − 10 f(x) = x2+ 6x + 8 f(x) = x2 − 6x − 8 f(x) = x2− 8

OpenStudy (jungleman03):

I think its a

OpenStudy (anonymous):

just square (x + 3) and then distribute the negative sign(-)

OpenStudy (anonymous):

\[-(x + 3)^2 − 1= -(x^2+6x+9)-1=-x^2-6x-9-1=-x^2-6x-10=\] \[-x^2-6x-10\]